OpenStudy (anonymous):

what is the solution to the inequality will fan and medal 7

OpenStudy (anonymous):

\[7\le b+8 and \frac{ b }{3} -2 \le-7?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, for the first inequality subtract 7 from sides. for the second inequality, first add 2 to both sides then multiply both sides by 3

OpenStudy (anonymous):

\[7leb+1

OpenStudy (anonymous):

sorry one sec

OpenStudy (anonymous):

right, now subtract one from both sides. you'll get \[-1 \le b \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for the second inequality add 2 to both sides and then multiply both sides by 3

OpenStudy (anonymous):

\[\frac{ 2b }{ 5 }\le-5\]

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

\[\frac{b}3-2+2\le -7 +2\]

OpenStudy (anonymous):

sorry im really bad at math

OpenStudy (anonymous):

\[\frac{b}{3} \le -5\] Practice makes perfect :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

now multiply both sides by 3 \[\frac{b}{3} \times 3 \le -5 \times3 \]

OpenStudy (anonymous):

\[b \le -15\] is the answer

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

You are welcome :)